OpenStudy (anonymous):

"A company produces CD players for a unit cost of $15.00 per CD player. The company has fixed costs of $120.00. If each CD player can be sold for $30.00, how many CD players must be sold to break even? Determine the cost equation first. Next, determine the revenue equation. Use the two equations you have found to determine the break-even point." Could somebody explain to me how to set up this system of linear equations? I figured the break even point would be when the fixed cost equals the profit made, so 120 = 30x - 15y where x is the amount of players sold, and y is the amount produced.

OpenStudy (anonymous):

\[15p+120=30p\], where p = the number of cd player sold.

OpenStudy (anonymous):

x=120 + 15c y=30 c x=y so 120 + 15c = 30c 120=15c 8=c you Can do it your way, but they break even point will be in an economical sense the absolute value of the y intercept